OpenStudy (anonymous):

The formula A = 200e0.033t models the population of a particular city, in thousands, t years after 1998. When will the population of the city reach 297 thousand? A) 2013 B) 2012 C) 2010 D) 2011

OpenStudy (anonymous):

@surdawi

OpenStudy (anonymous):

a=200e^0.033t so 200/297=.033

OpenStudy (anonymous):

write the equation with all the number

OpenStudy (anonymous):

ln(200/297)

OpenStudy (anonymous):

no :P

OpenStudy (anonymous):

relax, 1st write the equation will all the numbers

OpenStudy (anonymous):

well first the population before 1998 was 200

OpenStudy (anonymous):

297=200e^0.033t

OpenStudy (anonymous):

then

OpenStudy (anonymous):



OpenStudy (anonymous):

yes yes go

OpenStudy (anonymous):

ln(297/200)= 0.033t ln e

OpenStudy (anonymous):

ln(297/200)= 0.033t * 1

OpenStudy (anonymous):